OpenStudy (anonymous):

Simplify the expression the root of negative 1 all over the quantity of 2 minus 4 i minus the quantity of 5 plus 3 i.

OpenStudy (zehanz):

Do yo mean: \[\frac{ i }{ 2-4i-(5+3i) }\]

OpenStudy (anonymous):

@ZeHanz yes exactly what i mean

OpenStudy (zehanz):

Then the first step is to calculate the difference in the denominator:\[\frac{ i }{ (2-5) -(4+3)i }=\frac{ i }{ -3-7i }\]Next is a standard trick to get rid of the complex number in the denominator: multiply nominator and denominator with -3+7i (the conjugate of -3-7i):\[\frac{ i }{ -3-7i }\frac{ -3+7i }{ -3+7i }=\] Because you know what happens when you write out (a-b)(a+b), can you do the rest yourself now?

OpenStudy (anonymous):

so @ZeHanz whats the answer?

OpenStudy (anonymous):

@ZeHanz ???

OpenStudy (zehanz):

\[\frac{ i(-3+7i) }{ 9+49 }=\frac{ -7-3i }{ 58 }\]